Company issued 1,000 bonds each with a par value of $1,000 and each carrying a detachable warrant for the purchase of 4 shares of stock at $40 each. On the day of the bond issue, the bond could have been issued without warrants at par while the warrants could have been sold separately for $60. The amount received for the combined instrument on the date of the issue was $1,025,000. Using the proportional method, record the issue of the two securities.
Answer #1

A B С D $1,000 $60 $1,000,000 $60,000 0.9434 0.0566 1 2 Fair value of bond without detachable warrant 3 Fair value of detachaA B С 1000 60 1000000 60000 =B5/(B5+B6) =B6/(B5+B6) 1 2 Fair value of bond without detachable warrant 3 Fair value of detacha